OpenStudy (mollykenney):

Use substitution to solve each system of equations. Check your solution. x − 5y = 2 9x + 8 = 15y

Nnesha (nnesha):

so you should get one of equation for x or either for y

Nnesha (nnesha):

so i would recommend first one is easy to solve for x so we can substitution in another equation

Nnesha (nnesha):

can you solve first equation for x ??

OpenStudy (mollykenney):

9x+8 = 15 (5y + 2) ??

OpenStudy (nikato):

not quite. if you add 5y to both sides in the first equation, you get x=2+5y

OpenStudy (nikato):